A company has an overhead crane that has an estimated remaining life of 8 years. The crane can be sold now for $6000. If the crane is kept in service, it must be overhauled immediately at a cost of $3000. Operating and maintenance costs will be $2800 per year after the crane is overhauled. The overhauled crane will have a $500 MV at the end of the 8-year study period. A new crane will cost $22000, will last for 8 years, and will have a $2000 MV at that time. Operating and maintenance costs are $1000 per year for the new crane. The company uses a before-tax interest rate of 10% per year in evaluating investment alternatives.

a)What is the AW of the defender?

b)What is the AW of the challenger?

c)Should the company replace the old crane?
Answer #1

a) AW of the defender = -11000*(A/P,10%,8)-2800

= -11000*0.187444-2800

= -4861.88

b) AW of the challenger = -22000*(A/P,10%,8)-1000+2000*(A/F,10%,8)

= -22000*0.187444-1000+2000*0.087444

= -4948.88

c) The company should not replace the old crane